OpenStudy (anonymous):

What is equivalent to the expression sqr root of -121? A. 11i B. -11i C. sqr root 11i D. - sqr root 11 E. - sqr root 11i

OpenStudy (anonymous):

\[\sqrt{-121}= \sqrt{121\times-1}=11\times \sqrt{-1}=11i\]

OpenStudy (anonymous):

u find it good ?